# Wolfpack Adversarial Attack for Robust Multi-Agent Reinforcement Learning

### v2

@@ E.1 Comparison Results for Other CTDE Algorithms
Method Map Name
2s3z
$K_{\mathrm{WP}}=2$
$t_{\mathrm{WP}}=3$
3m
$K_{\mathrm{WP}}=1$
$t_{\mathrm{WP}}=3$
3s_vs_3z
$K_{\mathrm{WP}}=2$
$t_{\mathrm{WP}}=3$
8m
$K_{\mathrm{WP}}=1$
$t_{\mathrm{WP}}=3$
MMM
$K_{\mathrm{WP}}=2$
$t_{\mathrm{WP}}=3$
1c3s5z
$K_{\mathrm{WP}}=2$
$t_{\mathrm{WP}}=3$
Mean
Natural Vanilla QPLEX $97.1\pm 1.25$ $\mathbf{99.2\pm 0.45}$ $99.4\pm 0.50$ $97.4\pm 0.33$ $99.5\pm 0.47$ $97.1\pm 0.94$ $98.4\pm 0.07$
RANDOM $97.7\pm 2.11$ $98.7\pm 0.78$ $99.6\pm 0.09$ $99.1\pm 0.87$ $98.2\pm 1.23$ $98.5\pm 1.25$ $98.7\pm 0.54$
RARL $94.4\pm 4.54$ $89.7\pm 1.98$ $88.4\pm 6.80$ $97.8\pm 1.63$ $98.8\pm 0.82$ $94.6\pm 1.60$ $94.0\pm 1.35$
RAP $96.5\pm 1.70$ $96.6\pm 2.49$ $93.2\pm 0.90$ $99.1\pm 0.47$ $98.8\pm 0.78$ $97.8\pm 0.82$ $97.7\pm 0.39$
ERNIE $96.7\pm 2.54$ $98.8\pm 0.82$ $99.7\pm 0.08$ $98.8\pm 1.42$ $99.1\pm 0.50$ $98.4\pm 0.97$ $98.6\pm 0.55$
ROMANCE $97.1\pm 2.88$ $93.2\pm 5.88$ $98.8\pm 0.82$ $94.7\pm 3.18$ $\mathbf{99.7\pm 0.05}$ $99.0\pm 1.08$ $96.8\pm 1.19$
WALL (ours) $\mathbf{99.5\pm 0.52}$ $97.7\pm 2.13$ $\mathbf{99.9\pm 0.08}$ $\mathbf{99.8\pm 0.14}$ $99.0\pm 0.65$ $\mathbf{99.5\pm 0.64}$ $\mathbf{99.2\pm 0.29}$
Random Attack Vanilla QPLEX $75.2\pm 2.54$ $36.2\pm 6.63$ $81.9\pm 16.0$ $40.0\pm 9.83$ $65.3\pm 6.18$ $75.0\pm 2.79$ $62.3\pm 5.64$
RANDOM $85.9\pm 0.87$ $69.0\pm 5.77$ $96.2\pm 1.61$ $89.0\pm 7.67$ $91.2\pm 2.41$ $95.5\pm 1.19$ $87.8\pm 0.65$
RARL $81.2\pm 13.6$ $61.3\pm 13.9$ $74.7\pm 12.6$ $73.7\pm 21.6$ $92.0\pm 0.82$ $92.0\pm 2.94$ $79.1\pm 4.27$
RAP $90.8\pm 2.83$ $78.0\pm 6.98$ $92.1\pm 10.3$ $65.3\pm 4.50$ $95.0\pm 1.41$ $95.7\pm 1.89$ $85.0\pm 1.88$
ERNIE $82.0\pm 2.02$ $59.0\pm 8.98$ $93.0\pm 1.63$ $80.0\pm 1.41$ $91.3\pm 0.94$ $93.0\pm 0.82$ $83.1\pm 1.89$
ROMANCE $89.2\pm 2.05$ $64.9\pm 13.4$ $93.7\pm 2.05$ $51.6\pm 5.26$ $92.2\pm 4.12$ $95.2\pm 0.99$ $76.0\pm 2.84$
WALL (ours) $\mathbf{97.4\pm 0.86}$ $\mathbf{85.4\pm 2.45}$ $\mathbf{99.4\pm 0.45}$ $\mathbf{92.9\pm 3.49}$ $\mathbf{98.3\pm 1.16}$ $\mathbf{98.6\pm 1.23}$ $\mathbf{95.3\pm 0.69}$
EGA Vanilla QPLEX $48.1\pm 3.40$ $16.0\pm 5.11$ $72.5\pm 15.1$ $58.5\pm 16.8$ $71.0\pm 4.24$ $76.9\pm 1.46$ $57.2\pm 6.15$
RANDOM $60.5\pm 6.78$ $61.4\pm 14.9$ $82.0\pm 6.56$ $86.2\pm 3.39$ $85.3\pm 2.53$ $88.8\pm 2.53$ $77.4\pm 3.10$
RARL $63.4\pm 0.45$ $65.7\pm 6.24$ $71.0\pm 8.16$ $86.0\pm 6.53$ $89.7\pm 1.89$ $84.7\pm 2.05$ $76.7\pm 0.99$
RAP $78.5\pm 4.02$ $83.3\pm 0.94$ $85.4\pm 5.26$ $89.0\pm 2.16$ $92.7\pm 1.70$ $96.7\pm 0.47$ $88.0\pm 0.72$
ERNIE $64.4\pm 6.99$ $67.3\pm 10.3$ $51.7\pm 9.46$ $86.7\pm 4.71$ $86.0\pm 4.97$ $89.3\pm 4.50$ $74.2\pm 3.01$
ROMANCE $79.5\pm 6.02$ $80.3\pm 1.65$ $90.3\pm 1.25$ $80.7\pm 8.25$ $95.2\pm 2.02$ $92.5\pm 1.27$ $85.6\pm 1.04$
WALL (ours) $\mathbf{89.0\pm 3.77}$ $\mathbf{83.9\pm 2.88}$ $\mathbf{99.6\pm 0.57}$ $\mathbf{94.4\pm 1.23}$ $\mathbf{96.4\pm 1.18}$ $\mathbf{96.1\pm 0.70}$ $\mathbf{93.2\pm 0.70}$
Wolfpack Adversarial Attack (ours) Vanilla QPLEX $30.8\pm 4.32$ $11.8\pm 7.46$ $63.7\pm 24.3$ $30.7\pm 11.1$ $20.0\pm 12.5$ $41.9\pm 6.08$ $33.1\pm 7.08$
RANDOM $50.5\pm 2.92$ $16.8\pm 6.76$ $89.5\pm 5.70$ $45.6\pm 18.0$ $34.0\pm 14.1$ $82.5\pm 11.3$ $53.2\pm 4.91$
RARL $55.5\pm 2.05$ $27.3\pm 9.18$ $78.0\pm 4.32$ $46.0\pm 14.9$ $21.3\pm 14.4$ $79.7\pm 6.94$ $51.3\pm 2.81$
RAP $59.0\pm 4.72$ $50.3\pm 14.7$ $86.7\pm 8.23$ $33.7\pm 3.09$ $45.0\pm 6.98$ $93.7\pm 2.36$ $56.3\pm 3.24$
ERNIE $52.9\pm 6.28$ $49.0\pm 9.90$ $76.0\pm 9.90$ $42.7\pm 10.6$ $20.7\pm 10.4$ $78.7\pm 6.65$ $53.3\pm 3.56$
ROMANCE $57.3\pm 8.39$ $38.5\pm 13.9$ $89.7\pm 1.25$ $28.6\pm 1.10$ $46.2\pm 10.6$ $83.5\pm 5.53$ $50.8\pm 0.89$
WALL (ours) $\mathbf{88.3\pm 1.20}$ $\mathbf{87.6\pm 4.72}$ $\mathbf{99.7\pm 0.47}$ $\mathbf{84.5\pm 1.67}$ $\mathbf{96.3\pm 3.86}$ $\mathbf{99.3\pm 0.90}$ $\mathbf{92.6\pm 1.32}$
@@ E.3 Performance Comparison for EGA
This section presents the training performance of WALL under the existing Evolutionary Generation-based Attackers (EGA) (Yuan et al., 2023) method across six SMAC scenarios, as shown in Figure E.4. The results demonstrate that WALL consistently achieves superior robustness across all scenarios, even against unseen EGA adversaries that are not included in its training process.
@@ Appendix F Additional Ablation studies
Number of Wolfpack Attacks $K_{\mathrm{WP}}$: The hyperparameter $K_{\mathrm{WP}}$ determines the number of Wolfpack attacks, with each attack consisting of an initial attack and follow-up attacks over $t_{\mathrm{WP}}=3$ timesteps. The total number of attack steps $K$ for Wolfpack attack is then calculated as $K=4\times K_{\mathrm{WP}}$. In this section, we e conduct a parameter search for $K_{\mathrm{WP}}\in[1,2,3,4]$. Fig. F.1 illustrates the robustness of WALL policies trained with different $K_{\mathrm{WP}}$ values under the default Wolfpack attack in 8m and MMM. In both environments, having too small $K_{\mathrm{WP}}$ results in insufficiently severe attacks, which leads to reduced robustness of the WALL framework. Conversely, in the 8m environment, excessively large $K_{\mathrm{WP}}$ values create overly devastating attacks, making it difficult for CTDE methods to learn strategies to counter the Wolfpack attack, which degrades learning performance. Therefore, an optimal $K_{\mathrm{WP}}$ exists in both environments: $K_{\mathrm{WP}}=2$ for 8m and $K_{\mathrm{WP}}=4$ for MMM, which we choose as the default hyperparameters.
@@ G.1 Visualization of Wolfpack Adversarial Attack Across Additional SMAC Scenarios
Fig. G.2(b) illustrates Vanilla QMIX under the Wolfpack adversarial attack. An initial attack is launched at $t=6$, followed by the targeting of four follow-up agents between $t=7$ and $t=9$. The follow-up agents selected include one healing agent attempting to heal the initial agent, one guarding agent positioned to protect the initial agent, and two agents actively engaging the enemy targeting the initial agent. Due to the Wolfpack attack, the initial agent failed to receive critical healing or guarding support at $t=10$ and $t=19$, leading to its elimination. This disruption severely hindered the remaining agents’ ability to defend against the adversarial attack, ultimately resulting in a loss as all agents are defeated.
@@ G.2 Additional Analysis of Follow-up Agent Group Selection
Figure G.4: $Soft(Q^{i})$ and $Soft(\tilde{Q}^{i})$ for each agent, along with the difference between the two distributions
Figure G.5: KL divergence values for each agent, representing the difference between $Soft(Q^{i})$ and $Soft(\tilde{Q}^{i})$ distributions
Appendix H Limitations
The proposed WALL framework effectively defends against the highly detrimental Wolfpack adversarial attack, but it has certain limitations. One drawback is the reliance on complex Transformers for identifying critical steps, which adds computational overhead. However, our method is not strictly dependent on Transformers; any recurrent network capable of generating sequences can be employed in its place. Additionally, while parameter search is required to optimize the Wolfpack attack, the results demonstrate that our method consistently outperforms existing single-agent attack algorithms regardless of parameter choice. Addressing these limitations presents opportunities for further research and improvement.

### v3

@@ E.1 Comparison Results for Other CTDE Algorithms
Method Scenario 2s3z 3m 3s_vs_3z 8m MMM 1c3s5z Mean
Natural Vanilla QPLEX $97.1\pm 1.3$ $\mathbf{99.2\pm 0.5}$ $99.4\pm 0.5$ $97.4\pm 0.3$ $99.5\pm 0.5$ $97.1\pm 0.9$ $98.4\pm 0.1$
RANDOM $97.7\pm 2.1$ $98.7\pm 0.8$ $99.6\pm 0.1$ $99.1\pm 0.9$ $98.2\pm 1.2$ $98.5\pm 1.3$ $98.7\pm 0.5$
RARL $94.4\pm 4.5$ $89.7\pm 2.0$ $88.4\pm 6.8$ $97.8\pm 1.6$ $98.8\pm 0.8$ $94.6\pm 1.6$ $94.0\pm 1.4$
RAP $96.5\pm 1.7$ $96.6\pm 2.5$ $93.2\pm 0.9$ $99.1\pm 0.5$ $98.8\pm 0.8$ $97.8\pm 0.8$ $97.7\pm 0.4$
ERNIE $96.7\pm 2.5$ $98.8\pm 0.8$ $99.7\pm 0.1$ $98.8\pm 1.4$ $99.1\pm 0.5$ $98.4\pm 1.0$ $98.6\pm 0.6$
ROMANCE $97.1\pm 2.9$ $93.2\pm 5.9$ $98.8\pm 0.8$ $94.7\pm 3.2$ $\mathbf{99.7\pm 0.1}$ $99.0\pm 1.1$ $96.8\pm 1.2$
WALL (ours) $\mathbf{99.5\pm 0.5}$ $97.7\pm 2.1$ $\mathbf{99.9\pm 0.1}$ $\mathbf{99.8\pm 0.1}$ $99.0\pm 0.7$ $\mathbf{99.5\pm 0.6}$ $\mathbf{99.2\pm 0.3}$
Random Attack Vanilla QPLEX $75.2\pm 2.5$ $36.2\pm 6.6$ $81.9\pm 16.0$ $40.0\pm 9.8$ $65.3\pm 6.2$ $75.0\pm 2.8$ $62.3\pm 5.6$
RANDOM $85.9\pm 0.9$ $69.0\pm 5.8$ $96.2\pm 1.6$ $89.0\pm 7.7$ $91.2\pm 2.4$ $95.5\pm 1.2$ $87.8\pm 0.7$
RARL $81.2\pm 13.6$ $61.3\pm 13.9$ $74.7\pm 12.6$ $73.7\pm 21.6$ $92.0\pm 0.8$ $92.0\pm 2.9$ $79.1\pm 4.3$
RAP $90.8\pm 2.8$ $78.0\pm 7.0$ $92.1\pm 10.3$ $65.3\pm 4.5$ $95.0\pm 1.4$ $95.7\pm 1.9$ $85.0\pm 1.9$
ERNIE $82.0\pm 2.0$ $59.0\pm 9.0$ $93.0\pm 1.6$ $80.0\pm 1.4$ $91.3\pm 0.9$ $93.0\pm 0.8$ $83.1\pm 1.9$
ROMANCE $89.2\pm 2.1$ $64.9\pm 13.4$ $93.7\pm 2.1$ $51.6\pm 5.3$ $92.2\pm 4.1$ $95.2\pm 1.0$ $76.0\pm 2.8$
WALL (ours) $\mathbf{97.4\pm 0.9}$ $\mathbf{85.4\pm 2.5}$ $\mathbf{99.4\pm 0.5}$ $\mathbf{92.9\pm 3.5}$ $\mathbf{98.3\pm 1.2}$ $\mathbf{98.6\pm 1.2}$ $\mathbf{95.3\pm 0.7}$
EGA Vanilla QPLEX $48.1\pm 3.4$ $16.0\pm 5.1$ $72.5\pm 15.1$ $58.5\pm 16.8$ $71.0\pm 4.2$ $76.9\pm 1.5$ $57.2\pm 6.2$
RANDOM $60.5\pm 6.8$ $61.4\pm 14.9$ $82.0\pm 6.6$ $86.2\pm 3.4$ $85.3\pm 2.5$ $88.8\pm 2.5$ $77.4\pm 3.1$
RARL $63.4\pm 0.5$ $65.7\pm 6.2$ $71.0\pm 8.2$ $86.0\pm 6.5$ $89.7\pm 1.9$ $84.7\pm 2.1$ $76.7\pm 1.0$
RAP $78.5\pm 4.0$ $83.3\pm 0.9$ $85.4\pm 5.3$ $89.0\pm 2.2$ $92.7\pm 1.7$ $96.7\pm 0.5$ $88.0\pm 0.7$
ERNIE $64.4\pm 7.0$ $67.3\pm 10.3$ $51.7\pm 9.5$ $86.7\pm 4.7$ $86.0\pm 5.0$ $89.3\pm 4.5$ $74.2\pm 3.0$
ROMANCE $79.5\pm 6.0$ $80.3\pm 1.7$ $90.3\pm 1.3$ $80.7\pm 8.3$ $95.2\pm 2.0$ $92.5\pm 1.3$ $85.6\pm 1.0$
WALL (ours) $\mathbf{89.0\pm 3.8}$ $\mathbf{83.9\pm 2.9}$ $\mathbf{99.6\pm 0.6}$ $\mathbf{94.4\pm 1.2}$ $\mathbf{96.4\pm 1.2}$ $\mathbf{96.1\pm 0.7}$ $\mathbf{93.2\pm 0.7}$
Wolfpack Adversarial Attack (ours) Vanilla QPLEX $30.8\pm 4.3$ $11.8\pm 7.5$ $63.7\pm 24.3$ $30.7\pm 11.1$ $20.0\pm 12.5$ $41.9\pm 6.1$ $33.1\pm 7.1$
RANDOM $50.5\pm 2.9$ $16.8\pm 6.8$ $89.5\pm 5.7$ $45.6\pm 18.0$ $34.0\pm 14.1$ $82.5\pm 11.3$ $53.2\pm 4.9$
RARL $55.5\pm 2.1$ $27.3\pm 9.2$ $78.0\pm 4.3$ $46.0\pm 14.9$ $21.3\pm 14.4$ $79.7\pm 6.9$ $51.3\pm 2.8$
RAP $59.0\pm 4.7$ $50.3\pm 14.7$ $86.7\pm 8.2$ $33.7\pm 3.1$ $45.0\pm 7.0$ $93.7\pm 2.4$ $56.3\pm 3.2$
ERNIE $52.9\pm 6.3$ $49.0\pm 9.9$ $76.0\pm 9.9$ $42.7\pm 10.6$ $20.7\pm 10.4$ $78.7\pm 6.7$ $53.3\pm 3.6$
ROMANCE $57.3\pm 8.4$ $38.5\pm 13.9$ $89.7\pm 1.3$ $28.6\pm 1.1$ $46.2\pm 10.6$ $83.5\pm 5.5$ $50.8\pm 0.9$
WALL (ours) $\mathbf{88.3\pm 1.2}$ $\mathbf{87.6\pm 4.7}$ $\mathbf{99.7\pm 0.5}$ $\mathbf{84.5\pm 1.7}$ $\mathbf{96.3\pm 3.9}$ $\mathbf{99.3\pm 0.9}$ $\mathbf{92.6\pm 1.3}$
@@ E.3 Performance Comparison for EGA
This section presents the training performance of WALL under the existing Evolutionary Generation-based Attackers (EGA) (Yuan et al., 2023) method across six SMAC scenarios, as shown in Fig. E.4. The results demonstrate that WALL consistently achieves superior robustness across all scenarios, even against unseen EGA adversaries that are not included in its training process.
@@
E.4 Computational Cost
We measured the total training time (for 3M steps) required for each algorithm in the SMAC environments. The results are summarized in Table E.3. WALL requires approximately 30% more time than ROMANCE, primarily due to the use of the Transformer. However, despite this additional cost, the Transformer-based critical step selection plays an essential role in the effectiveness of our method, as it enables precise identification of attack timing where coordinated perturbations are most disruptive. This capability allows the robust policy to anticipate and defend against strategically timed adversarial threats, ultimately resulting in stronger robustness. In general, training on the MMM map takes longer than on the 8m map due to the increased number of agents, which leads to higher computational costs as a result of scalability.
Method Scenario 8m MMM
[…]
Table E.3: Training time comparison on 8m and MMM
E.5 General Robustness Experiments
We conducted additional experiments to evaluate robustness under other commonly used criteria in the 8m and MMM environments. Specifically, we considered: (1) Gaussian observation noise: Standard Gaussian noise (mean 0, standard deviation 1) is injected into agents’ local observations, where attack steps (total attack steps : 8 for 8m, 16 for MMM) and the attack group are selected randomly; and (2) Different parameterization in the SMAC test environment: Robustness is evaluated under perturbed unit attributes by reducing the initial health of allied units by 10%, 15%, and 20% in the test setup, compared to the training configuration. These types of noise and perturbations are standard in robustness evaluations and enable us to assess how well the proposed method generalizes beyond action-level perturbations. The Table E.4 below summarizes the results. In both settings, WALL demonstrates consistently stronger performance than existing baselines, suggesting that it generalizes well to broader forms of distributional shift.
Method Scenario 8m MMM Mean
Gaussian Obs. Noise Vanilla QMIX $62.6\pm 2.3$ $75.3\pm 3.1$ $69.0\pm 2.1$
RANDOM $72.3\pm 4.2$ $79.3\pm 3.9$ $75.8\pm 7.2$
ROMANCE $69.6\pm 11.2$ $76.6\pm 7.5$ $73.1\pm 6.6$
WALL (ours) $\mathbf{91.3\pm 3.3}$ $\mathbf{97.3\pm 1.7}$ $\mathbf{94.3\pm 3.3}$
Ally HP ↓ 10% Vanilla QMIX $52.8\pm 4.8$ $88.4\pm 2.7$ $70.6\pm 1.8$
RANDOM $50.8\pm 8.2$ $95.4\pm 2.5$ $73.1\pm 5.2$
ROMANCE $56.8\pm 16.0$ $92.4\pm 4.5$ $74.6\pm 8.5$
WALL (ours) $\mathbf{73.2\pm 7.5}$ $\mathbf{98.6\pm 1.1}$ $\mathbf{85.9\pm 3.7}$
Ally HP ↓ 15% Vanilla QMIX $47.4\pm 6.9$ $69.0\pm 5.4$ $58.2\pm 5.8$
RANDOM $49.2\pm 4.5$ $81.4\pm 5.5$ $65.3\pm 4.6$
ROMANCE $57.6\pm 15.7$ $81.2\pm 3.5$ $69.4\pm 7.8$
WALL (ours) $\mathbf{69.2\pm 10.5}$ $\mathbf{94.0\pm 3.2}$ $\mathbf{81.6\pm 5.7}$
Ally HP ↓ 20% Vanilla QMIX $0.3\pm 0.4$ $41.2\pm 4.5$ $20.8\pm 2.1$
RANDOM $0.0\pm 0$ 0.0 $65.0\pm 2.7$ $32.5\pm 1.3$
ROMANCE $2.3\pm 0.4$ $57.0\pm 8.9$ $29.7\pm 4.3$
WALL (ours) $\mathbf{4.3\pm 1.9}$ $\mathbf{89.6\pm 5.6}$ $\mathbf{47.0\pm 3.3}$
Table E.4: Average test win rates of robust MARL policies under various perturbation settings
@@ Appendix F Additional Ablation studies
Number of Wolfpack Attacks $K_{\mathrm{WP}}$: The hyperparameter $K_{\mathrm{WP}}$ determines the number of Wolfpack attacks, with each attack consisting of an initial attack and follow-up attacks over $t_{\mathrm{WP}}=3$ timesteps. The total number of attack steps $K$ for Wolfpack attack is then calculated as $K=4\times K_{\mathrm{WP}}$. In this section, we conduct a parameter search for $K_{\mathrm{WP}}\in[1,2,3,4]$. Fig. F.1 illustrates the robustness of WALL policies trained with different $K_{\mathrm{WP}}$ values under the default Wolfpack attack in 8m and MMM. In both environments, having too small $K_{\mathrm{WP}}$ results in insufficiently severe attacks, which leads to reduced robustness of the WALL framework. Conversely, in the 8m environment, excessively large $K_{\mathrm{WP}}$ values create overly devastating attacks, making it difficult for CTDE methods to learn strategies to counter the Wolfpack attack, which degrades learning performance. Therefore, an optimal $K_{\mathrm{WP}}$ exists in both environments: $K_{\mathrm{WP}}=2$ for 8m and $K_{\mathrm{WP}}=4$ for MMM, which we choose as the default hyperparameters.
@@ G.1 Visualization of Wolfpack Adversarial Attack Across Additional SMAC Scenarios
Fig. G.2(b) illustrates Vanilla QMIX under the Wolfpack adversarial attack. An initial attack is launched at $t=6$, followed by the targeting of four follow-up agents between $t=7$ and $t=9$. The follow-up agents selected include one healing agent attempting to heal the initial agent, one guarding agent positioned to protect the initial agent, and two agents actively engaging the enemy targeting the initial agent. Due to the Wolfpack attack, the initial agent failed to receive critical healing or guarding support at $t=10$ and $t=19$, leading to its elimination. This disruption severely hinders the remaining agents’ ability to defend against the adversarial attack, ultimately resulting in a loss as all agents are defeated.
@@ G.2 Additional Analysis of Follow-up Agent Group Selection
Figure G.4: $Soft(Q^{i})$ and $Soft(\tilde{Q}^{i})$ for each agent, along with the difference between the two distributions.
Figure G.5: KL divergence values for each agent, representing the difference between $Soft(Q^{i})$ and $Soft(\tilde{Q}^{i})$ distributions.